OpenStudy (anonymous):

Can anyone help me on this problem? Provide a written explanation describing a step by step process to find the vertical asymptotes of a rational function.

OpenStudy (anonymous):

Set the denominator of the rational function equal to 0. This is where the function does not exist, a vertical asymptote.

OpenStudy (anonymous):

Phydaeux is right. Sometimes it's notated as y=/= x or as VA @ x

OpenStudy (anonymous):

\[f(x)=(x^3-2x)/(2(x^2-5))\] 2x^2-10 =0 2x^2=10 x^2=5 \[x=\sqrt{5}\] So you have a vertical asymptote at \[x=\sqrt{5}\]